### v1

# An Innovative Attention-based Ensemble System for Credit Card Fraud Detection

Abstract
Detecting credit card fraud (CCF) holds significant importance due to its role in safeguarding consumers from unauthorized transactions that have the potential to result in financial detriment and negative impacts on their credit rating. It aids financial institutions in upholding the reliability of their payment mechanisms and circumventing the expensive procedure of compensating for deceitful transactions. The utilization of Artificial Intelligence methodologies demonstrated remarkable efficacy in the identification of credit card fraud instances. Within this study, we present a unique attention-based ensemble model. This model is enhanced by adding an attention layer for integration of first layer classifiers’ predictions and a selection layer for choosing the best integrated value. The attention layer is implemented with two aggregation operators: dependent ordered weighted averaging (DOWA) and induced ordered weighted averaging (IOWA). The performance of the IOWA operator is very close to the learning algorithm in neural networks which is based on the gradient descent optimization method, and performing the DOWA operator is based on weakening the classifiers that make outlier predictions compared to other learners. Both operators have a sufficient level of complexity for the recognition of complex patterns. Accuracy and diversity are the two criteria we use for selecting the classifiers whose predictions are to be integrated by the two aggregation operators. Using a bootstrap forest, we identify the 13 most significant features of the dataset that contribute the most to CCF detection and use them to feed the proposed model. Exhibiting its efficacy, the ensemble model attains an accuracy of 99.95% with an area under the curve (AUC) of 1.
Index Terms:
Credit card fraud detection, Fusion-based parallel ensemble system, Attention layer, Machine learning, Classifier selection, Aggregation operators.
I Introduction
The convenience of financial services for consumers around the world has been tremendously improved by digital technology. However, this development also opened up new avenues for financial fraud, especially in credit card transactions [1, 2], [3]. Criminals defraud people through various technologies and false identities intending to steal their money [4]. This gave way to a steep increase in credit card frauds as e-commerce and online shopping gained momentum [5, 6].
Looking at the numbers, one can tell that credit card fraud (CCF) is a big concern. According to Nilson’s report; global losses from card fraud amounted to 28.65 billion dollars in 2019. These figures are expected surpassing 43 billion dollars by 2025 [7, 8]. The large increase shows that fraudulent methods are getting more complicated day by day, hence making it hard to beat fraudsters at their game [9, 10]. In the United States alone, CCF cases reported to the Federal Trade Commission (FTC) were over 393 thousand during 2020, indicating a significant rise from the previous year [11, 12]. The advent of internet-based transactions worsened matters leading into rise of fraud cases involving card-not-present (CNP) transactions, which are quite common currently [1, 13]. Compared to point-of-sale fraud, the prevalence of CNP fraud rose by 81%, suggesting a major shift in digital crime [14]. This therefore means that it is necessary to propose models and techniques for detecting deceitful transactions. The old ways of identifying fraud were not quick enough [15, 16]. As a result, researchers performed thorough investigations and practically used machine learning (ML) algorithms specifically designed for CCF detection [17, 18, 19, 20].
Fraud detection extensively relies on the utilization of ML algorithms. Various ML algorithms, like support vector machine (SVM), decision tree, random forest, LightGBM, XGBoost, naïve bayes, stacking classifier and voting classifier are utilized in fraud detection [21, 22, 23, 21, 24].
Deep learning (DL) techniques were widely used in detecting fraud across fields in CCF detection. Models like Convolutional Neural Networks (CNN), Recurrent Neural Networks, and Autoencoders displayed results in spotting activities [25, 1, 26]. Hybrid models such as CNN-SVM, Autoencoder-LightGBM, ensemble method and others were commonly employed for CCF detection [27, 28, 29, 30].
Our key contributions to this research are:
  • We introduce a novel fusion-based ensemble system by adding an attention layer and a selection module to the traditional stacking classifier.
  • Intelligently, two aggregation operators, DOWA and IOWA, are employed to combine the predictions of first layer classifiers. In this entirely fusion-based architecture, a synergistic combination of the aggregated values by these two operators is used to feed the meta-learner.
  • Dimension reduction using the bootstrap forest approach is accomplished by eliminating features that contribute negligibly to the prediction accuracy.
  • The best classifiers are selected for combination based on two criteria: performance and diversity.
  • The proposed model demonstrates 99.95% accuracy in detecting credit card fraud with the AUC of 1.
The rest of this paper is structured as follows: Section II introduces related work on detecting CCF. Section III first addresses the dataset and the identification of its most significant features, then introduces the aggregation operators used in the attention layer, and finally analyses the structure of the proposed system. Findings and analysis are examined in Section IV. Section V presents prospects for future studies. Finally, Section VI considers the conclusion.
II Background and Related Work
Several methodologies and algorithms were deployed to predict fraud, which saw a sharp rise in recent times. A focal area of various studies is the detection of CCF. These paragraphs delve into some of these research efforts.
Long et al. A graph network model was introduced to efficiently detect online fraud by addressing challenges such as feature disparities, category imbalances, and relationship issues [31, 32, 33, 34].
In [12], the authors proposed a CCF detection model using new transactional behavioral representations. Two time-aware gates were designed in a recurrent neural network to capture long- and short-term transactional habits and behavioral changes influenced by time intervals between transactions. The model used a time-aware attention module to extract behavioral information from historical transactions, considering time intervals, and an interaction module to enhance the learned representations. Similarly, in the realm of spotting news, LSTM and Bi LSTM methods were utilized for text data preprocessing and analysis, achieving a level of accuracy in detecting misinformation [35, 36].
Fidelis and coworkers in [37] presented a novel approach using a hybrid cultural genetic algorithm (CGA) alongside a trained modular neural network ensemble for classifying transactions. This ensemble included a complex architecture of modular neural networks, featuring components such as an encoder, selectors, a swapped recombine, swapped dynamic amplification, and a belief finisher for CGA operations. The ensemble relied on optimized rules derived from the GA-block for effective classification tasks.
Ibomoiye et al. in [38] introduced a DL ensemble method to identify CCF utilizing a mix of LSTM and gated recurrent unit (GRU) networks, as primary learners alongside an MLP acting as the meta-learner. The combined model demonstrated superior performance in sensitivity, specificity and AUC metrics when contrasted with performing GRU, LSTM and MLP models.
Within [39], the authors presented an innovative approach to detecting fraud, comprising four distinct stages. Initially, cardholders were classified into groups based on similar transaction behaviors using historical transaction data. Next, a window-sliding strategy is utilized to merge transactions in each group. Following that, distinct behavioral patterns are derived from every cardholder. Separate classifiers are trained for each group, enabling real-time fraud detection.
In [40], the authors propose the utilization of an adaptive sampling and aggregation-based graph neural network (ASA-GNN) in order to enhance transaction fraud detection. By employing a neighbor sampling strategy that selectively filters out noisy nodes and provides supplementary information for fraudulent nodes, the model effectively addresses the challenge of learning discriminative representations. To selectively choose similar neighbors with comparable behavior patterns, cosine similarity and edge weights are utilized, while introducing a neighbor diversity metric to address fraudsters’ attempts at disguise and prevent excessive smoothing.
In [41], Wang and colleagues introduced an ensemble approach that incorporates various feature ranking methods to achieve more reliable results. This method involved a two-step process: first, employing multiple rankers to create different feature rankings, and then combining these rankings using a median-based method to determine the top ten features for model building. This approach led to the development of models utilizing Decision Tree-based classifiers, specifically CatBoost and XGBoost, enhancing predictive stability and dependability.
Esenogho et al. described an ensemble classifier combining an LSTM network with the AdaBoost algorithm for efficient credit card fraud detection in [42]. Here, the LSTM serves as the base learner within AdaBoost, using data refined by the synthetic minority over-sampling technique-edited nearest neighbor (SMOTE-ENN) technique for better model training. The ensemble’s final decision was reached through a weighted voting scheme among the outcomes of various LSTM models, aiming to boost fraud detection accuracy.
In [43], the authors introduced a fresh approach to detect consumer fraud in reviews. Their method involves integrating a channel biattention convolutional neural network (CNN) with a pre-trained language model. A novel similarity computation module is introduced, which employs a metric matrix to evaluate the connection between prior knowledge and consumer reviews. This enables the model to effectively identify previously undetected fraudulent behaviors.
Although various efforts were made towards the detection and prediction of CCF, many focused on single classifiers, which are not as reliable as ensemble-based classification methods. This is because multi-classifier approaches use the consensus of several single classifiers, resulting in a more reliable outcome. Studies that examined CCF detection using ensemble classifiers either employed outdated versions such as bagging, boosting, and stacking.
Figure 1: Heatmap of correlation matrix.
III Proposed Attention-based Ensemble System for CCF Detection
In this section, a new parallel ensemble architecture based on information fusion is introduced. This novel architecture is constructed by adding two layers to the stacking classifier structure: a) an attention layer, and b) a selection layer. Information fusion in the attention layer is accomplished through two operators, DOWA and IOWA, while the selection layer employs a mathematical logic to choose one of the two output vectors from the attention layer for each sample. Ultimately, the selected vector feeds the meta-learner.
III-A Dataset
The dataset in question includes over 550,000 anonymized records of credit card transactions performed by European cardholders in 2023 [44]. The anonymization was meticulously done to ensure the privacy and security of the cardholders’ identities. This dataset’s main purpose is to aid in creating and honing fraud detection algorithms and systems capable of pinpointing transactions that might be fraudulent. Some principal characteristics of this dataset are:
  • The dataset comprises a unique identifier, named “id” assigned to each transaction for distinct identification.
  • It includes anonymized features labeled V1 through V28 that encapsulate varying aspects of the transactions, such as timing and geographical details, among others, in a manner protective of user privacy.
  • “Amount” denotes the monetary value of the transaction.
  • there is a binary attribute named “Class” which serves the purpose of classifying transactions into two categories: fraudulent (marked as 1) and legitimate (marked as 0).
Figure 2: this boxplot visualizes the distribution of dataset features along with outliers.
Fig. 1 shows the correlation matrix, which displays the extent of linear dependencies among different features of the dataset through a heatmap. The more a cell corresponding to two features tends towards violet or white, it indicates higher linear dependencies. The boxplot in Fig. 2 Offers a graphical representation of the essential characteristics of the dataset, encompassing the median, upper and lower quartiles, and potential outliers.
Figure 3: Predictor screening report for feature selection
III-B Predictor screening
The process of predictor screening is vital for singling out significant predictors from a broad pool of potential ones. This technique uses a bootstrap forest method to identify predictors that could impact the outcome variable.
In the analysis, a set of 100 decision trees is created using the bootstrap forest approach for each response. This model evaluates how each predictor influences the prediction of outcome. The predictors are ranked based on their contributions to the model. Fig. 3 illustrates this ranking, displaying the predictors with their corresponding contributions.
Hence, predictors whose contribution percentage falls below 0.5% are classified as insignificant dimensions. By removing these negligible dimensions from feature space $R^{29}$, we effectively transition into feature space $R^{13}$. Thus, as per Fig. 1 and Fig. 3, 16 features that have high linear or non-linear dependency on each other or are not sufficiently discriminative are removed from the raw dataset to prevent the model from becoming biased towards these uninformative features.
III-C aggregation operators
The OWA operator is a mathematical tool used for merging inputs by considering both their order and assigned weights, frequently applied in processes of decision-making and aggregation. This method systematically orders the inputs before applying specific weights to them, as highlighted in [45]. Definition 1 introduces the DOWA operator, and the IOWA operator is introduced in definition 2. These two operators are used separately to combine the predictions of first layer classifiers in subsection III-D.
Definition 1. Let $p_{1},p_{2},...,p_{n}$ represent the probabilistic outputs or forecasts from initial classifiers regarding a particular sample, indicating the likelihood of the sample’s classification into each category. Let $m$ denote the mean of these predictions across all classes, i.e., $m=\frac{1}{n}\displaystyle\sum\limits_{j=1}^{n}p_{j},\ (\alpha(1),\alpha(2),...,\alpha(n))$ is a permutation of $(1,2,...,n)$ such that $p_{\alpha(j-1)}\geq p_{\alpha(j)}$ for all $j=2,...,n$, then we call
$c(p_{\alpha(j)},m)=1-\frac{|p_{\alpha(j)}-m|}{\sum\limits_{j=1}^{n}|p_{j}-m|},\ \ \ \ j=1,2,...,n,$ (1)
the similarity degree between the $j$-th highest value $p_{\alpha(j)}$ and the mean value $m$.
The DOWA operator calculates the weighted mean of the predictions using a set of weights, each of which emphasizes the relevance or significance of the respective prediction. Denote by $w=(w_{1},w_{2},...,w_{n})^{T}$ the weight vector for the OWA operator, and we establish:
$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{\alpha(j)},m)},\ \ \ \ j=1,2,...,n,$ (2)
where $c(p_{\alpha(j)},m)$ is defined by (1). Clearly, we have $w_{j}\in[0,1]$ and $\sum\limits_{j=1}^{n}w_{j}=1$. Since
$\sum\limits_{j=1}^{n}c(p_{\alpha(j)},m)=\sum\limits_{j=1}^{n}c(p_{j},m),$ (3)
then (2) can be reformulated as
$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{j},m)}.\ \ \ \ j=1,2,...,n,$ (4)
Under these circumstances, it follows that
$OWA(p_{1},p_{2},...,p_{n})=\frac{\sum\limits_{j=1}^{n}c(p_{\alpha(j)},m)p_{\alpha(j)}}{\displaystyle\sum\limits_{j=1}^{n}c(p_{j},m)}.$ (5)
Using (4) weights corresponding to the predictions of each classifier are calculated. These weights can have different values for each sample. Then, the predictions of the first layer classifiers are aggregated based on the weights assigned to them by (5). This operator effectively converts the separate and independent predictions of single first layer classifiers into a single prediction.
Definition 2. The development of the weighting vector is a critical aspect widely discussed, as seen in [46, 47]. Filev and Yager proposed a method in [48] for deriving the weighting vector for an OWA aggregation through the analysis of observational data. This approach closely aligns with learning algorithms typical of neural networks, as referenced in [49, 50], utilizing the gradient descent method as its foundation. This section outlines the process for determining the OWA weighting vector’s weights based on observational input.
We are presented with a set of samples, each comprising a series of values $(p_{k1},p_{k2},...,p_{kn})$, referred to as the arguments, and a corresponding single value named the aggregated value, denoted as $d_{k}$. Our aim is to develop a model that accurately captures this aggregation process using the OWA (Ordered Weighted Averaging) method. In essence, the task is to derive a weighting vector $W$ that effectively represents the aggregation mechanism across the dataset. The aim is to determine the weighting vector $W$ based on the data provided.
The task focuses on simplifying the challenge of identifying the weights by leveraging the OWA aggregation’s linear nature with the reordered inputs. For the kth data point, the reordered arguments are signified as $b_{k1},b_{k2},...,b_{kn}$, with $b_{kj}$ being the jth largest value within the set $(p_{k1},p_{k2},...,p_{kn})$. With the arguments now ordered, the objective turns into determining the OWA weight vector $W^{T}=[w_{1},w_{2},...,w_{n}]$ in such a manner that the equation $b_{k1}w_{1}+b_{k2}w_{2}+...+b_{kn}w_{n}=d_{k}$ holds true for each instance k, from 1 up to N.
We will adjust our method by searching for a weights vector, $W$, for OWA, aiming to closely resemble the aggregating function by reducing immediate errors $e_{k}$
$e_{k}=\frac{1}{2}\ (b_{k1}w_{1}+b_{k2}w_{2}+...+b_{kn}w_{n}-d_{k})^{2}.$ (6)
In relation to the weights $w_{i}$, this learning challenge represents a constrained optimization issue. This is because the OWA weights $w_{i}$ must adhere to two specific criteria: $\sum\limits_{i=1}^{n}w_{i}=1$ and $w_{i}\in[0,1]$ for $i=1$ to $n$.
To bypass the restrictions, each of the OWA weights is represented in the following manner.
$w_{i}=\frac{e^{\beta{i}}}{\sum\limits_{j=1}^{n}e^{\beta{j}}}.$ (7)
With the transformation, regardless of the parameter values $\beta_{i}$, the weights $w_{i}$ will fall within the unit interval and their sum will equal 1. Hence, the original constrained minimization issue is converted into an unconstrained nonlinear programming problem:
Minimize the immediate errors $\textbf{e}_{k}:$
By integrating (7) into (6), it can be stated:
$e_{k}=\frac{1}{2}\ (b_{k1}\frac{e^{\beta_{1}}}{\sum\limits_{j=1}^{n}e^{\beta_{j}}}+b_{k2}\frac{e^{\beta_{2}}}{\sum\limits_{j=1}^{n}e^{\beta_{j}}}+...+b_{kn}\frac{e^{\beta_{n}}}{\sum\limits_{j=1}^{n}e^{\beta_{j}}}-d_{k})^{2}$
Motivated by the significant achievements of gradient descent methodologies in the backpropagation approach utilized for training in neural networks, we adopt it here. By applying the gradient descent technique, we establish the subsequent rule for parameter updates $\beta_{i},i=(1,n)$;
$\beta_{i}(l+1)=\beta_{i}(l)-\alpha\frac{\partial e_{k}}{\partial\beta_{i}}|_{\beta_{i}=\beta_{i}(l)},$ (9)
where $\alpha$ represents the rate of learning $(0<\alpha<1)$ and $\beta_{i}(l)$ signifies the approximation of $\beta_{i}$ following the lth iteration.
To simplify notation, we represent the estimate of the aggregated value $d_{k}$ as $\hat{d}_{k}$
$\hat{d}_{k}=b_{k1}w_{1}+b_{k2}w_{2}+...+b_{kn}w_{n}.$ (10)
Then, regarding the partial derivative $\frac{\partial e_{k}}{\partial\beta_{i}}$ we get
$\frac{\partial e_{k}}{\partial\beta_{i}}=w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k}),\ \ \ i=[1,n].$ (11)
Finally, we establish the formula for adjusting the parameters $\beta_{i}$ in the following manner:
$\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k})$ (12)
where the $w_{i}$ values are determined at each iterative step based on the current estimates of the $\beta_{i}$ values
$w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}},\ \ i=[1,n].$ (13)
and $\hat{d}_{k}$ represents the current approximation of the aggregated values $d_{k}$.
This outlines the procedure implemented at each step of the iteration.
  • We possess a current estimation of the $\beta_{i}$, denoted as $\beta_{i}(l)$, and a new observation composed of the ordered arguments $b_{k1},b_{k2},...,b_{kn}$ along with an aggregated value $d_{k}$.
  • We utilize $\beta_{i}(l)$ to generate a current estimation for the weights
    $w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}}.$ (14)
  • We apply the estimated weights and the ordered arguments to derive a computed aggregated value
    $\hat{d}_{k}=b_{k1}w_{1}(l)+b_{k2}w_{2}(l)+...+b_{kn}w_{n}(l).$ (15)
  • We update our estimations of the $\lambda_{i}$.
    $\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(l)(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k}).$ (16)
Using the 10-fold cross-validation method, we divide the predicted probabilities into two groups: test data and training data. Then, with the help of the proposed algorithm, we learn the weights from the training data, and then, using these weights and the sorted input arguments, the aggregation value is calculated for the test data.
III-D CCF detection Framework: Structure, Development, and Examination
Figure 4: Traditional stacking classifier
Figure 5: The block diagram of proposed architecture for detecting CCF.
Fig. 4 illustrates the structure of the traditional stacking classifier. Stacking is a method of ensemble learning that merges predictions from base models to generate a final prediction.
These base models can be of various types, including decision trees, SVM, neural networks, etc. The predictions made by these base models are then used as input features for a second-level model (or meta-learner) that aims to correct the predictions of the base models and produce the final prediction.
This study enhances the traditional stacking architecture by incorporating an attention layer and a selection layer. Fig. 5 shows the block diagram of the proposed architecture for CCF detection. In the following, we delve into the details of this entirely attention-based parallel architecture.
III-D1 First layer classifier selection
Selecting first layer classifiers is a crucial step in building an effective stacking model. The diversity and performance of these classifiers have a significant impact on the stacking model’s overall performance. Here are key considerations and strategies for selecting first layer classifiers (base models) in a stacking classifier:
Performance
Using feeble models can harm the overall performance, as they may introduce too much noise. However, models do not need to be top performers individually, as their errors might be corrected by the meta-learner. The best candidates for first layer classifiers are those not only with the best performance individually, but also whose predictions possess an acceptable level of diversity. Under these circumstances, the classifiers will complement each other well. Table I shows the 6 base classifiers that were able to meet the above criteria.
Diversity
Diverse base classifiers are crucial for making errors in different parts of the dataset. Similar classifiers lead to similar mistakes, reducing stacking effectiveness. Correlation between the predictions of first layer classifiers is a measure for assessing the level of diversity. Fig. 6 shows the correlation values between the predictions for each pair of first-level classifiers. It highlights how these two classifiers’ predictions are connected or related. This could be visualized using lines, colors, or patterns indicating areas, with medium or low correlation. High correlation areas indicate that both classifiers typically agree on sample classification, while low correlation areas suggest disagreement.
Figure 6: The correlation among the predictions generated by the classifiers in the first layer.
III-D2 attention layer
At the heart of the suggested design lies the attention layer, which allows the model to concentrate on classifiers. This layer allows the model to assign separate weightings to the predictions of each classifier based on the belief associated with each classifier. Considering the DOWA aggregation operator, which assigns a negligible weight to the classifier prediction that is an outlier compared to the other two classifiers, it is necessary for the classifiers, whose predictions are to be fused by this operator, to have sufficiently close correlations in their predictions. For example, if we assume the correlation between classifiers $C1$, $C2$, and $C3$ to be $C1C2:0.85$, $C1C3:0.9$, and $C2C3:0.4$, then by aggregating the predictions of these three classifiers using the DOWA operator, this operator will disregard the impact of one of the two classifiers, either $C2$ or $C3$, for each sample. Therefore, to employ this operator in its best performance, we select the classifiers Boosted Trees (BT), Extra Trees (ET), and AdaBoost (AB), which according to Fig. 6 have prediction correlations of 0.69, 0.64, and 0.56, respectively, which are significantly close to each other, for aggregation using DOWA. The predictions of the remaining three classifiers, namely SGD, SVM, and MLP, are aggregated using the IOWA operator.
Therefore, the weights are calculated for the three classifiers BT, ET, and AB by the DOWA aggregation operator, and for the three classifiers SGD, SVM, and MLP by the IOWA operator. Ultimately, the fusion of the probabilities assigned to each sample by the first layer classifiers is done using the calculated weights, and these probabilities are sorted in descending order. The output of the DOWA operator for each sample $i=1,2,...,N$ is a vector in the form of
$F_{DOWA}=[f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0}),f(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})]^{T},$ (17)
and similarly, the output of the IOWA operator is a vector in the form of
$\small F_{IOWA}=[f(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0}),f(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})]^{T},$ (18)
where $p_{BT}^{0}$indicates the probability of sample i belonging to class 0 as classified by BT, and $f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$ represents the fused value of absolute probabilities assigned to class zero by the three classifiers BT, ET, and AB. Considering that the DOWA and IOWA aggregation operators each aggregate the soft outputs of the three specified classifiers, it can be said that each of these operators is a mapping from the space $R^{6}$ to $R^{2}$.
III-D3 selection layer
This layer receives two vectors, $F_{DOWA}$ and $F_{IOWA}$, as inputs. The basis of its operation for each sample is according to the Algorithm 1.
Algorithm 1 : Sample-Based Aggregated Value Selector
Input: $f_{i}(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$, $f_{i}(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})$, $f_{i}(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0})$, $f_{i}(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})$
Output: One of the aggregate values, $F_{DOWA}$ or $F_{IOWA}$, for each sample
$i\leftarrow 1$
while $i\leq N$ do
     if $|f_{i}(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})-f_{i}(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})|\geq|f_{i}(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0})-f_{i}(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})|$ then
         return $F_{DOWA}$
     else
         return $F_{IOWA}$
     end if
     $i\leftarrow i+1$
end while
In fact, for each sample, this layer decides to intelligently select one of the two vectors, $F_{DOWA}$ or $F_{IOWA}$. The logic of selection is that whichever of the two operators, DOWA or IOWA, has greater confidence in a sample belonging to one of the two classes, 0 or 1, the aggregated value by that operator is chosen. In this way, we use a smart selection of aggregated values by both operators to feed the meta-learner. It can be said that the selection layer synergistically combines the results from these two operators.
III-D4 Meta-learner selection
The stacking process is significantly influenced by a meta-learner, as it plays a crucial role in determining the classifier’s performance. The selection of a classifier for meta learning is crucial as it directly influences the stacked ensemble’s generalization on unseen instances and, subsequently, its overall predictive accuracy. In the context of the stacking architecture, the Ridge classifier can be considered a favorable option due to its moderate complexity, which facilitates interpretability. Additionally, it incorporates regularization techniques to mitigate overfitting and effectively handles potential multicollinearity among the features used for learning. In Table II, the test results are presented, indicating that the Ridge classifier outperforms other base classifiers.
TABLE I: The degree of accuracy achieved by the first layer classifiers when evaluated on the test data.
Classifier Accuracy
Boosted Trees (BT) $93.64\pm 0.22\%$
Stochastic Gradient Descent (SGD) $92.48\pm 0.13\%$
Extra Trees (ET) $94.09\pm 0.31\%$
AdaBoost (AB) $95.33\pm 0.19\%$
Support vector Machine (SVM) $94.77\pm 0.25\%$
Multi-Layer Perceptron (MLP) $95.87\pm 0.23\%$
TABLE II: Accuracy metric for choosing the meta-learner.
Learner Accuracy
Decision Tree $95.32\pm 0.16\%$
Gradient Boosting $96.71\pm 0.13\%$
K-nearest Neighbor $96.83\pm 0.11\%$
Logistic Regression $98.44\pm 0.19\%$
Ridge Classifier $99.95\pm 0.04\%$
CatBoost $97.63\pm 0.19\%$
IV Findings and Analysis
In this segment, we delve deeper into the results and present a range of performance indicators to evaluate the effectiveness of the proposed algorithm. These include measures such as accuracy, sensitivity (also known as recall), precision, specificity, Matthew’s Correlation Coefficient (MCC), and the F1-Score.
IV-A Performance metrics
IV-A1 Precision
Precision quantifies the ratio of correct positive predictions to the total number of positive predictions made by a model, highlighting the accuracy of its positive predictions.
$Precision=\frac{TP}{TP+FP}.$ (19)
IV-A2 Specificity
Specificity measures the ratio of accurate negative predictions to the total negative predictions made by a model, underlining the precision of its negative predictions, e.g.,
$Specificity=\frac{TN}{TN+FP}.$ (20)
IV-A3 Accuracy
Accuracy assesses the total proportion of both true positive and true negative predictions made by a model compared to the overall number of predictions, reflecting the general reliability of the model’s predictions, e.g.,
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}.$ (21)
IV-A4 Sensitivity
Sensitivity quantifies the percentage of true positive instances accurately identified by a model, emphasizing the model’s capacity to detect all positive cases, e.g.,
$Sensitivity=\frac{TP}{TP+FN}.$ (22)
IV-A5 MCC
Unlike other metrics that may only focus on the positive class or are overly optimistic in imbalanced datasets, MCC provides a balanced measure, even if the classes are of very different sizes. MCC is defined mathematically as:
$\small MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}.$ (23)
IV-A6 F1-Score
The F1-score is a harmonized measure that considers precision and recall. It offers a singular metric that equilibrates the balance between precision and recall, proving valuable in situations involving datasets with imbalances, e.g.,
$F1-Score=\frac{2\times Precision\times sensitivity}{Precision+sensitivity},$ (24)
where $TP$ is true positive, $FP$ is false positive, $TN$ is true negative and $FN$ is false negative.
To prevent data leakage at all stages, including predictions by first layer classifiers, getting weights through a learning process by the IOWA operator, and the final prediction by the meta-learner, the 10-fold cross-validation method was used. The exceptional performance of the proposed architecture on the test data aligns with the findings reported in Table III.
Fig. 7 presents the receiver operating characteristic (ROC) curve, which showcases the correlation between the true positive rate (TPR) and the false positive rate (FPR) across different classification thresholds. The ROC curves for both classes demonstrate a significant level of overlap, suggesting comparable performance between the two classes. The ROC curve’s area under the curve (AUC) functions as a metric to assess classifier performance. A higher AUC value, approaching 1, signifies superior classifier performance. The suggested model demonstrates the AUC value of 1.
Figure 7: ROC curve along with corresponding AUC on test data.
TABLE III: Performance metrics of the proposed method.
Metric Value
Precision $100\%$
Accuracy $99.95\%$
Sensitivity $99.90\%$
Specificity $100\%$
MCC $99.91\%$
F1-Score $99.94\%$
V Prospects for Future Studies
In future research, we aim to explore:
  • In the proposed architecture, we randomly selected 6 classifiers as the first layer classifiers. Future work could optimize the number of first layer classifiers by relying on forward selection and backward elimination approaches.
  • Future studies are recommended to focus on other aggregation methods, such as fuzzy integral operators and approaches with a higher level of complexity than OWA operators.
  • There are various criteria for selecting first layer classifiers that future work could employ and compare.
  • Upcoming studies could implement the suggested architecture for additional datasets that differ in terms of data origin and class balance and evaluate its performance.
VI Conclusion
The importance of detecting CCF lies in its ability to protect financial assets, maintain consumer confidence, ensure regulatory compliance, and enhance the overall security and integrity of the financial system. In this study, we introduced a novel architecture that is entirely based on integration for the detection of CCF. This architecture combines two traditional stacking architectures in parallel through the aid of two key layers including attention and selection. The attention layer utilizes two aggregation operators, DOWA and IOWA, to decide how to combine the first layer classifiers’ predictions. Each aggregation operator is tasked with combining predictions of three distinct classes. The selection layer decides which one of the two operators, DOWA or IOWA, should be considered for feeding the meta-learner for each sample, depending on which operator has more confidence in assigning the sample to one of the two classes, zero or one. This architecture, based on combining multiple classifiers, has led to improved overall accuracy, reliability, and decision-making process. This technique leverages the strengths of different classifiers and mitigates their individual weaknesses, potentially leading to more robust and accurate predictions, especially in complex or challenging classification tasks.

### v2

@@
-An Innovative Attention-based Ensemble System for Credit Card Fraud Detection
+Explainable AI for Fraud Detection: An Attention-Based Ensemble of CNNs, GNNs, and A Confidence-Driven Gating Mechanism
 Abstract
-Detecting credit card fraud (CCF) holds significant importance due to its role in safeguarding consumers from unauthorized transactions that have the potential to result in financial detriment and negative impacts on their credit rating.
-It aids financial institutions in upholding the reliability of their payment mechanisms and circumventing the expensive procedure of compensating for deceitful transactions.
-The utilization of Artificial Intelligence methodologies demonstrated remarkable efficacy in the identification of credit card fraud instances.
-Within this study, we present a unique attention-based ensemble model.
-This model is enhanced by adding an attention layer for integration of first layer classifiers’ predictions and a selection layer for choosing the best integrated value.
-The attention layer is implemented with two aggregation operators: dependent ordered weighted averaging (DOWA) and induced ordered weighted averaging (IOWA).
-The performance of the IOWA operator is very close to the learning algorithm in neural networks which is based on the gradient descent optimization method, and performing the DOWA operator is based on weakening the classifiers that make outlier predictions compared to other learners.
-Both operators have a sufficient level of complexity for the recognition of complex patterns.
-Accuracy and diversity are the two criteria we use for selecting the classifiers whose predictions are to be integrated by the two aggregation operators.
-Using a bootstrap forest, we identify the 13 most significant features of the dataset that contribute the most to CCF detection and use them to feed the proposed model.
-Exhibiting its efficacy, the ensemble model attains an accuracy of 99.95% with an area under the curve (AUC) of 1.
+The rapid expansion of e-commerce and the widespread use of credit cards in online purchases and financial transactions have significantly heightened the importance of promptly and accurately detecting credit card fraud (CCF).
+Not only do fraudulent activities in financial transactions lead to substantial monetary losses for banks and financial institutions, but they also undermine user trust in digital services.
+This study presents a new stacking-based approach for CCF detection by adding two extra layers to the usual classification process: an attention layer and a confidence-based combination layer.
+In the attention layer, we combine soft outputs from a convolutional neural network (CNN) and a recurrent neural network (RNN) using the dependent ordered weighted averaging (DOWA) operator, and from a graph neural network (GNN) and a long short-term memory (LSTM) network using the induced ordered weighted averaging (IOWA) operator.
+These weighted outputs capture different predictive signals, increasing the model’s accuracy.
+Next, in the confidence-based layer, we select whichever aggregate (DOWA or IOWA) shows lower uncertainty to feed into a meta-learner.
+To make the model more explainable, we use shapley additive explanations (SHAP) to identify the top ten most important features for distinguishing between fraud and normal transactions.
+These features are then used in our attention-based model.
+Experiments on three datasets show that our method achieves high accuracy and robust generalization, making it effective for CCF detection.
 Index Terms:
-Credit card fraud detection, Fusion-based parallel ensemble system, Attention layer, Machine learning, Classifier selection, Aggregation operators.
+Credit card fraud detection, Attention mechanism, Explainable AI, Information fusion, Ensemble learning.
 I Introduction
 The convenience of financial services for consumers around the world has been tremendously improved by digital technology.
-However, this development also opened up new avenues for financial fraud, especially in credit card transactions [1, 2], [3].
-Criminals defraud people through various technologies and false identities intending to steal their money [4].
-This gave way to a steep increase in credit card frauds as e-commerce and online shopping gained momentum [5, 6].
+However, this development also opened up new avenues for financial fraud, especially in credit card transactions [1, 2].
 Looking at the numbers, one can tell that credit card fraud (CCF) is a big concern.
 According to Nilson’s report; global losses from card fraud amounted to 28.65 billion dollars in 2019.
-These figures are expected surpassing 43 billion dollars by 2025 [7, 8].
-The large increase shows that fraudulent methods are getting more complicated day by day, hence making it hard to beat fraudsters at their game [9, 10].
+These figures are expected surpassing 38.5 billion dollars by 2027 [3, 4].
 In the United States alone, CCF cases reported to the federal trade commission (FTC) were over 393 thousand during 2020, indicating a significant rise from the previous year [5, 6].
 The advent of internet-based transactions worsened matters leading into rise of fraud cases involving card-not-present (CNP) transactions, which are quite common currently [1, 7].
-Compared to point-of-sale fraud, the prevalence of CNP fraud rose by 81%, suggesting a major shift in digital crime [14].
-This therefore means that it is necessary to propose models and techniques for detecting deceitful transactions.
-The old ways of identifying fraud were not quick enough [15, 16].
-As a result, researchers performed thorough investigations and practically used machine learning (ML) algorithms specifically designed for CCF detection [17, 18, 19, 20].
-Fraud detection extensively relies on the utilization of ML algorithms.
-Various ML algorithms, like support vector machine (SVM), decision tree, random forest, LightGBM, XGBoost, naïve bayes, stacking classifier and voting classifier are utilized in fraud detection [21, 22, 23, 21, 24].
-Deep learning (DL) techniques were widely used in detecting fraud across fields in CCF detection.
-Models like Convolutional Neural Networks (CNN), Recurrent Neural Networks, and Autoencoders displayed results in spotting activities [25, 1, 26].
-Hybrid models such as CNN-SVM, Autoencoder-LightGBM, ensemble method and others were commonly employed for CCF detection [27, 28, 29, 30].
-Our key contributions to this research are:
-  • We introduce a novel fusion-based ensemble system by adding an attention layer and a selection module to the traditional stacking classifier.
-  • Intelligently, two aggregation operators, DOWA and IOWA, are employed to combine the predictions of first layer classifiers.
-In this entirely fusion-based architecture, a synergistic combination of the aggregated values by these two operators is used to feed the meta-learner.
-  • Dimension reduction using the bootstrap forest approach is accomplished by eliminating features that contribute negligibly to the prediction accuracy.
-  • The best classifiers are selected for combination based on two criteria: performance and diversity.
-  • The proposed model demonstrates 99.95% accuracy in detecting credit card fraud with the AUC of 1.
-The rest of this paper is structured as follows: Section II introduces related work on detecting CCF.
-Section III first addresses the dataset and the identification of its most significant features, then introduces the aggregation operators used in the attention layer, and finally analyses the structure of the proposed system.
-Findings and analysis are examined in Section IV.
-Section V presents prospects for future studies.
-Finally, Section VI considers the conclusion.
-II Background and Related Work
-Several methodologies and algorithms were deployed to predict fraud, which saw a sharp rise in recent times.
-A focal area of various studies is the detection of CCF.
-These paragraphs delve into some of these research efforts.
-Long et al. A graph network model was introduced to efficiently detect online fraud by addressing challenges such as feature disparities, category imbalances, and relationship issues [31, 32, 33, 34].
-In [12], the authors proposed a CCF detection model using new transactional behavioral representations.
-Two time-aware gates were designed in a recurrent neural network to capture long- and short-term transactional habits and behavioral changes influenced by time intervals between transactions.
-The model used a time-aware attention module to extract behavioral information from historical transactions, considering time intervals, and an interaction module to enhance the learned representations.
-Similarly, in the realm of spotting news, LSTM and Bi LSTM methods were utilized for text data preprocessing and analysis, achieving a level of accuracy in detecting misinformation [35, 36].
-Fidelis and coworkers in [37] presented a novel approach using a hybrid cultural genetic algorithm (CGA) alongside a trained modular neural network ensemble for classifying transactions.
-This ensemble included a complex architecture of modular neural networks, featuring components such as an encoder, selectors, a swapped recombine, swapped dynamic amplification, and a belief finisher for CGA operations.
-The ensemble relied on optimized rules derived from the GA-block for effective classification tasks.
-Ibomoiye et al.
-in [38] introduced a DL ensemble method to identify CCF utilizing a mix of LSTM and gated recurrent unit (GRU) networks, as primary learners alongside an MLP acting as the meta-learner.
-The combined model demonstrated superior performance in sensitivity, specificity and AUC metrics when contrasted with performing GRU, LSTM and MLP models.
-Within [39], the authors presented an innovative approach to detecting fraud, comprising four distinct stages.
-Initially, cardholders were classified into groups based on similar transaction behaviors using historical transaction data.
-Next, a window-sliding strategy is utilized to merge transactions in each group.
-Following that, distinct behavioral patterns are derived from every cardholder.
-Separate classifiers are trained for each group, enabling real-time fraud detection.
-In [40], the authors propose the utilization of an adaptive sampling and aggregation-based graph neural network (ASA-GNN) in order to enhance transaction fraud detection.
-By employing a neighbor sampling strategy that selectively filters out noisy nodes and provides supplementary information for fraudulent nodes, the model effectively addresses the challenge of learning discriminative representations.
-To selectively choose similar neighbors with comparable behavior patterns, cosine similarity and edge weights are utilized, while introducing a neighbor diversity metric to address fraudsters’ attempts at disguise and prevent excessive smoothing.
-In [41], Wang and colleagues introduced an ensemble approach that incorporates various feature ranking methods to achieve more reliable results.
-This method involved a two-step process: first, employing multiple rankers to create different feature rankings, and then combining these rankings using a median-based method to determine the top ten features for model building.
-This approach led to the development of models utilizing Decision Tree-based classifiers, specifically CatBoost and XGBoost, enhancing predictive stability and dependability.
-Esenogho et al.
-described an ensemble classifier combining an LSTM network with the AdaBoost algorithm for efficient credit card fraud detection in [42].
-Here, the LSTM serves as the base learner within AdaBoost, using data refined by the synthetic minority over-sampling technique-edited nearest neighbor (SMOTE-ENN) technique for better model training.
-The ensemble’s final decision was reached through a weighted voting scheme among the outcomes of various LSTM models, aiming to boost fraud detection accuracy.
-In [43], the authors introduced a fresh approach to detect consumer fraud in reviews.
-Their method involves integrating a channel biattention convolutional neural network (CNN) with a pre-trained language model.
-A novel similarity computation module is introduced, which employs a metric matrix to evaluate the connection between prior knowledge and consumer reviews.
-This enables the model to effectively identify previously undetected fraudulent behaviors.
-Although various efforts were made towards the detection and prediction of CCF, many focused on single classifiers, which are not as reliable as ensemble-based classification methods.
-This is because multi-classifier approaches use the consensus of several single classifiers, resulting in a more reliable outcome.
-Studies that examined CCF detection using ensemble classifiers either employed outdated versions such as bagging, boosting, and stacking.
-Figure 1: Heatmap of correlation matrix.
-III Proposed Attention-based Ensemble System for CCF Detection
-In this section, a new parallel ensemble architecture based on information fusion is introduced.
-This novel architecture is constructed by adding two layers to the stacking classifier structure: a) an attention layer, and b) a selection layer.
-Information fusion in the attention layer is accomplished through two operators, DOWA and IOWA, while the selection layer employs a mathematical logic to choose one of the two output vectors from the attention layer for each sample.
-Ultimately, the selected vector feeds the meta-learner.
-III-A Dataset
-The dataset in question includes over 550,000 anonymized records of credit card transactions performed by European cardholders in 2023 [44].
-The anonymization was meticulously done to ensure the privacy and security of the cardholders’ identities.
-This dataset’s main purpose is to aid in creating and honing fraud detection algorithms and systems capable of pinpointing transactions that might be fraudulent.
-Some principal characteristics of this dataset are:
-  • The dataset comprises a unique identifier, named “id” assigned to each transaction for distinct identification.
-  • It includes anonymized features labeled V1 through V28 that encapsulate varying aspects of the transactions, such as timing and geographical details, among others, in a manner protective of user privacy.
-  • “Amount” denotes the monetary value of the transaction.
-  • there is a binary attribute named “Class” which serves the purpose of classifying transactions into two categories: fraudulent (marked as 1) and legitimate (marked as 0).
-Figure 2: this boxplot visualizes the distribution of dataset features along with outliers.
-Fig.
-1 shows the correlation matrix, which displays the extent of linear dependencies among different features of the dataset through a heatmap.
-The more a cell corresponding to two features tends towards violet or white, it indicates higher linear dependencies.
-The boxplot in Fig.
-2 Offers a graphical representation of the essential characteristics of the dataset, encompassing the median, upper and lower quartiles, and potential outliers.
-Figure 3: Predictor screening report for feature selection
-III-B Predictor screening
-The process of predictor screening is vital for singling out significant predictors from a broad pool of potential ones.
-This technique uses a bootstrap forest method to identify predictors that could impact the outcome variable.
-In the analysis, a set of 100 decision trees is created using the bootstrap forest approach for each response.
-This model evaluates how each predictor influences the prediction of outcome.
-The predictors are ranked based on their contributions to the model.
-Fig.
-3 illustrates this ranking, displaying the predictors with their corresponding contributions.
-Hence, predictors whose contribution percentage falls below 0.5% are classified as insignificant dimensions.
-By removing these negligible dimensions from feature space $R^{29}$, we effectively transition into feature space $R^{13}$.
-Thus, as per Fig.
-1 and Fig.
-3, 16 features that have high linear or non-linear dependency on each other or are not sufficiently discriminative are removed from the raw dataset to prevent the model from becoming biased towards these uninformative features.
-III-C aggregation operators
-The OWA operator is a mathematical tool used for merging inputs by considering both their order and assigned weights, frequently applied in processes of decision-making and aggregation.
-This method systematically orders the inputs before applying specific weights to them, as highlighted in [45].
-Definition 1 introduces the DOWA operator, and the IOWA operator is introduced in definition 2.
-These two operators are used separately to combine the predictions of first layer classifiers in subsection III-D.
+Traditional fraud detection methods, such as rule-based systems and manual transaction reviews, are often inadequate in the face of evolving fraud patterns [8, 9].
+These approaches are limited in their ability to adapt to new tactics employed by fraudsters and frequently generate high false-positive rates, leading to unnecessary inconvenience for legitimate customers [10, 11].
+In contrast, machine learning (ML) and deep learning (DL) methodologies offer data-driven solutions that are capable of learning complex patterns and adapting to emerging fraud strategies [12, 13].
+By analyzing vast amounts of transaction data, these techniques provide more accurate, scalable, and efficient ways to identify fraudulent transactions.
+Recent progress in data-driven computational intelligence has introduced powerful classification algorithms, such as decision trees, random forests, gradient boosting, convolutional neural networks (CNNs), and recurrent neural networks (RNNs), specifically tailored for fraud detection [14, 15, 16, 17].
+Additionally, ensemble techniques that synthesize the outputs of multiple models have shown notable success by exploiting the strengths of diverse classifiers [14, 18].
+However, despite these achievements, critical challenges persist in balancing predictive accuracy ($A_{c}$) with interpretability, particularly within financial institutions that require clear, trustworthy insights into automated decisions [19].
+Despite these notable advances, significant research gaps persist in CCF detection.
+First, many approaches still rely on a single classification model or a simple ensemble of models, which may fail to leverage complementary predictive signals and overlook more sophisticated aggregation strategies [20].
+Second, although graph neural networks (GNNs), CNNs, and recurrent architectures like RNNs and long short-term memories (LSTMs) each show promise individually, their combined outputs are often integrated through naive averaging or straightforward stacking, diluting critical predictive cues.
+Third, most current work does not adequately address model uncertainty, potentially leading to unreliable predictions, particularly in highly imbalanced datasets.
+Finally, interpretability remains a pressing concern: despite its importance in promoting stakeholder trust, the majority of methods do not offer transparent or user-friendly ways to understand feature importance and decision rationale.
+These limitations underscore the need for a more robust, uncertainty-aware, and interpretable framework that intelligently fuses multiple neural models to enhance both $A_{c}$ and trust in credit card fraud detection systems.
+Our main contributions to this study are outlined below:
+  • A diversity-based classifier selection is proposed, where an attention-based stacking framework is designed to leverage four distinct classifiers- CNN, RNN, LSTM, and GNN- to achieve a balanced trade-off between reliability and robustness.
+Specifically, CNN and RNN outputs are combined using the dependent ordered weighted averaging (DOWA) operator, while LSTM and GNN outputs are fused via the induced ordered weighted averaging (IOWA) operator.
+  • A confidence-aware combination layer is introduced, enabling dynamic selection of the more reliable aggregate (DOWA or IOWA) based on uncertainty estimates.
+This best-performing output is then fed into a meta-learner, ensuring higher $A_{c}$, particularly under highly imbalanced conditions.
+  • Interpretability is enhanced through the use of Shapley Additive Explanations (SHAP), allowing the identification of the top ten most influential features affecting model decisions.
+This feature-level insight improves transparency and facilitates stakeholder trust in the fraud detection process.
+  • An extensive evaluation and benchmarking are conducted using three benchmark datasets, where the proposed method is compared against individual classifiers (CNN, RNN, LSTM, GNN).
+The results demonstrate superior performance and robust generalization, validating the effectiveness of the proposed framework in credit card fraud detection.
+The remainder of this paper is organized as follows.
+Section II introduces the datasets used in our experiments and identifies the key features driving the proposed approach.
+It then discusses the aggregation operators (DOWA and IOWA) and concludes with a detailed exposition of the system’s architecture.
+Section III presents the experimental results, offering a comprehensive analysis of the findings.
+Section IV provides a broader discussion of these results, highlighting their implications and suggesting potential avenues for future research.
+Finally, Section V provides the concluding remarks.
+II Proposed Attention-based Ensemble System for CCF Detection
+In this section, we detail the methodology adopted to develop and evaluate our stacking-based approach for CCF detection.
+We begin by introducing the three benchmark datasets, highlighting their characteristics and the rationale behind their selection.
+We then describe the process for identifying key features using SHAP, focusing on the top attributes that distinguish fraudulent from legitimate transactions.
+Next, we explain how the DOWA and IOWA operators are employed to fuse the outputs of our neural network classifiers, CNN, RNN, LSTM, and GNN, in an attention mechanism.
+Finally, we outline the overall system architecture, including the confidence-based combination layer, which helps select the most reliable ensemble output before passing it to a meta-learner, ensuring both high $A_{c}$ and robust generalization.
+II-A Datasets
+For fraud detection, we use three datasets, each containing anonymized transaction-related features labeled $V1$ through $V28$ to preserve user privacy.
+These features capture various aspects of the transaction process, ranging from timing information to geographical context.
+Each dataset also includes the feature Amount, representing the monetary value of the transaction, and the binary attribute Class, where transactions are labeled as fraudulent (1) or legitimate (0).
+Table I provides a statistical overview of all three datasets.
+Notably, Dataset 3 was constructed by concatenating Dataset 1 and Dataset 2, resulting in a larger sample size for more comprehensive analysis and model training.
+To rigorously assess our framework in both balanced and imbalanced scenarios, we first employ a balanced dataset (Dataset 1) for initial training and testing.
+This approach allows us to observe the model’s baseline performance when both classes are represented more evenly.
+Next, we train our proposed framework on the highly imbalanced Dataset 2 and subsequently evaluate it on this same dataset to investigate how the skew in class distribution influences predictive $A_{c}$.
+Finally, we tune the model’s hyperparameters on the balanced dataset and test the refined model on the unbalanced dataset, mirroring real-world conditions where fraudulent transactions constitute only a small fraction of all transactions.
+Since the features in two datasets were anonymized, we assumed their equivalence when training on Dataset 1 and testing on Dataset 2.
+This step-by-step procedure ensures that our model is both reliable under balanced conditions and robust to the challenges posed by heavily imbalanced data.
+TABLE I: Data partition.
+Datasets samples Fraud Imbalance ratio
+Public 1 Train set 454 905 227 452 50%
+Test set 113 725 56 863 50%
+Public 2 Train set 227 845 394 0.17%
+Test set 56 962 98 0.17%
+Public 3 Train set 568 630 284 315 50%
+Test set 284 807 492 0.17%
+Figure 1: SHAP-based feature importance ranking for fraud detection.
+II-B Feature Importance Analysis Using SHAP
+We employ SHAP with a bootstrap-based approach using 100 decision trees to assess feature importance in our fraud detection model.
+The results, visualized in Fig.
+1, rank features by their average SHAP values.
+The analysis highlights $V14$, $V10$, and $V4$ as the most influential features, while others, such as $V24$ and $V27$, have minimal impact.
+Interestingly, the $Amount$ feature contributes negligibly, indicating that fraud detection relies more on transaction behavior than monetary values.
+To improve efficiency and reduce complexity, we selected only the top 10 features ($V14$, $V10$, $V4$, $V17$, $V12$, $V11$, $V16$, $V8$, $V3$, and $V1$) for the final model.
+This selection enhances interpretability, reduces overfitting risks, and ensures practical deployment while maintaining high predictive performance.
+II-C Aggregation Operators, DOWA and IOWA
+The Ordered Weighted Averaging (OWA) operator is a mathematical tool commonly employed in decision-making and aggregation tasks.
+It arranges inputs in a specified order before applying designated weights, as highlighted in [21].
+In this study, Definition 1 introduces the DOWA operator, while Definition 2 presents the IOWA operator.
+In section II-D, each operator is used independently to merge the predictions of the first-layer classifiers within our proposed framework.
 Definition 1.
@@
 then (2) can be reformulated as
-$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{j},m)}.\ \ \ \ j=1,2,...,n,$ (4)
+$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{j},m)},\ \ \ \ j=1,2,...,n,$ (4)
 under these circumstances, it follows that
@@
 We are presented with a set of samples, each comprising a series of values $(p_{k1},p_{k2},...,p_{kn})$, referred to as the arguments, and a corresponding single value named the aggregated value, denoted as $d_{k}$.
-Our aim is to develop a model that accurately captures this aggregation process using the OWA (Ordered Weighted Averaging) method.
+Our aim is to develop a model that accurately captures this aggregation process using the OWA method.
 In essence, the task is to derive a weighting vector $W$ that effectively represents the aggregation mechanism across the dataset.
@@
 Hence, the original constrained minimization issue is converted into an unconstrained nonlinear programming problem:
-Minimize the immediate errors $\textbf{e}_{k}:$
+Minimize the immediate errors $\mathbf{e_{k}}$ :
 By integrating (7) into (6), it can be stated:
@@
 Finally, we establish the formula for adjusting the parameters $\beta_{i}$ in the following manner:
-$\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k})$ (12)
+$\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k}),$ (12)
 where the $w_{i}$ values are determined at each iterative step based on the current estimates of the $\beta_{i}$ values
-$w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}},\ \ i=[1,n].$ (13)
+$w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}},\ \ i=[1,n],$ (13)
 and $\hat{d}_{k}$ represents the current approximation of the aggregated values $d_{k}$.
@@
     $\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(l)(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k}).$ (16)
-Using the 10-fold cross-validation method, we divide the predicted probabilities into two groups: test data and training data.
-Then, with the help of the proposed algorithm, we learn the weights from the training data, and then, using these weights and the sorted input arguments, the aggregation value is calculated for the test data.
-III-D CCF detection Framework: Structure, Development, and Examination
+In summary, both the DOWA and IOWA operators offer a systematic approach to aggregating probabilistic outputs by considering the order and relative similarity of predictions.
+By learning appropriate weights from training data and then applying them to sorted inputs, these operators capture subtle variations in classifier outputs, thereby enhancing the overall ensemble’s reliability.
+In the following section, we detail how these aggregation techniques are integrated into our proposed fraud detection framework, illustrating their role in the model’s multi-layer architecture.
+II-D CCF detection Framework: Structure, Development, and Examination
 Figure 2: Traditional stacking classifier
 Figure 3: The block diagram of proposed architecture for detecting CCF.
-Fig. 4 illustrates the structure of the traditional stacking classifier. Stacking is a method of ensemble learning that merges predictions from base models to generate a final prediction.
-These base models can be of various types, including decision trees, SVM, neural networks, etc.
-The predictions made by these base models are then used as input features for a second-level model (or meta-learner) that aims to correct the predictions of the base models and produce the final prediction.
-This study enhances the traditional stacking architecture by incorporating an attention layer and a selection layer.
+Fig.
+2 illustrates the structure of the traditional stacking classifier, an ensemble learning technique that improves predictive $A_{c}$ by combining multiple ML models.
+In this method, several learners are trained on the dataset, and their predictions are then used as input features for a meta-learner.
+The meta-learner learns how to optimally combine the outputs of the first layer models to make the final prediction.
+This study enhances the traditional stacking architecture by incorporating an attention layer and a confidence-aware combination layer.
 Fig.
 3 shows the block diagram of the proposed architecture for CCF detection.
-In the following, we delve into the details of this entirely attention-based parallel architecture.
-III-D1 First layer classifier selection
-Selecting first layer classifiers is a crucial step in building an effective stacking model.
-The diversity and performance of these classifiers have a significant impact on the stacking model’s overall performance.
-Here are key considerations and strategies for selecting first layer classifiers (base models) in a stacking classifier:
-Performance
-Using feeble models can harm the overall performance, as they may introduce too much noise.
-However, models do not need to be top performers individually, as their errors might be corrected by the meta-learner.
-The best candidates for first layer classifiers are those not only with the best performance individually, but also whose predictions possess an acceptable level of diversity.
-Under these circumstances, the classifiers will complement each other well.
-Table I shows the 6 base classifiers that were able to meet the above criteria.
-Diversity
-Diverse base classifiers are crucial for making errors in different parts of the dataset.
-Similar classifiers lead to similar mistakes, reducing stacking effectiveness.
-Correlation between the predictions of first layer classifiers is a measure for assessing the level of diversity.
-Fig.
-6 shows the correlation values between the predictions for each pair of first-level classifiers.
-It highlights how these two classifiers’ predictions are connected or related.
-This could be visualized using lines, colors, or patterns indicating areas, with medium or low correlation.
-High correlation areas indicate that both classifiers typically agree on sample classification, while low correlation areas suggest disagreement.
-Figure 6: The correlation among the predictions generated by the classifiers in the first layer.
-III-D2 attention layer
-At the heart of the suggested design lies the attention layer, which allows the model to concentrate on classifiers.
-This layer allows the model to assign separate weightings to the predictions of each classifier based on the belief associated with each classifier.
-Considering the DOWA aggregation operator, which assigns a negligible weight to the classifier prediction that is an outlier compared to the other two classifiers, it is necessary for the classifiers, whose predictions are to be fused by this operator, to have sufficiently close correlations in their predictions.
-For example, if we assume the correlation between classifiers $C1$, $C2$, and $C3$ to be $C1C2:0.85$, $C1C3:0.9$, and $C2C3:0.4$, then by aggregating the predictions of these three classifiers using the DOWA operator, this operator will disregard the impact of one of the two classifiers, either $C2$ or $C3$, for each sample.
-Therefore, to employ this operator in its best performance, we select the classifiers Boosted Trees (BT), Extra Trees (ET), and AdaBoost (AB), which according to Fig.
-6 have prediction correlations of 0.69, 0.64, and 0.56, respectively, which are significantly close to each other, for aggregation using DOWA.
-The predictions of the remaining three classifiers, namely SGD, SVM, and MLP, are aggregated using the IOWA operator.
-Therefore, the weights are calculated for the three classifiers BT, ET, and AB by the DOWA aggregation operator, and for the three classifiers SGD, SVM, and MLP by the IOWA operator.
-Ultimately, the fusion of the probabilities assigned to each sample by the first layer classifiers is done using the calculated weights, and these probabilities are sorted in descending order.
-The output of the DOWA operator for each sample $i=1,2,...,N$ is a vector in the form of
-$F_{DOWA}=[f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0}),f(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})]^{T},$ (17)
-and similarly, the output of the IOWA operator is a vector in the form of
-$\small F_{IOWA}=[f(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0}),f(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})]^{T},$ (18)
-where $p_{BT}^{0}$indicates the probability of sample i belonging to class 0 as classified by BT, and $f(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$ represents the fused value of absolute probabilities assigned to class zero by the three classifiers BT, ET, and AB.
-Considering that the DOWA and IOWA aggregation operators each aggregate the soft outputs of the three specified classifiers, it can be said that each of these operators is a mapping from the space $R^{6}$ to $R^{2}$.
-III-D3 selection layer
-This layer receives two vectors, $F_{DOWA}$ and $F_{IOWA}$, as inputs. The basis of its operation for each sample is according to the Algorithm 1.
-Algorithm 1 : Sample-Based Aggregated Value Selector
-Input: $f_{i}(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})$, $f_{i}(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})$, $f_{i}(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0})$, $f_{i}(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})$
+In the following, we delve into the details of this entirely attention-based ensemble architecture.
+II-D1 Diversity-Based Selection Strategy
+To evaluate the performance of deep learning models in fraud detection, we compare four architectures: CNN, RNN, LSTM, and GNN.
+These models are trained and optimized using cross-validation to ensure fair comparison.
+The main hyperparameters for each method are as follows.
+(1) CNN [27]: Number of convolutional layers = 3, Kernel size = 3, Stride = 1, Pooling type = MaxPooling, Pool size = 2, Activation function = ReLU, Fully connected layers = 2 (with 128 and 64 neurons), Optimizer = Adam (learning rate = 0.001), Epochs = 200.
+(2) RNN [28]: Number of recurrent layers = 2, Hidden units per layer = 128, Activation function = Tanh, Dropout rate = 0.2, Optimizer = RMSprop (learning rate = 0.001), Epochs = 200,
+(3) LSTM [28]: Number of LSTM layers = 2, Hidden units per layer = 128, Activation function = Tanh, Dropout rate = 0.3, Recurrent dropout = 0.2, Optimizer = Adam (learning rate = 0.0005), Epochs = 200.
+(4) GNN [29]: Number of graph convolution layers = 3, Hidden units per layer = 64, Graph aggregation method = Mean pooling, Activation function = LeakyReLU, Dropout rate = 0.2, Optimizer = Adam (learning rate = 0.001), Epochs = 200.
+Based on the correlation matrix in Fig.
+4, we observe that RNN and LSTM have the highest correlation (0.87), indicating considerable overlap in their predictive patterns.
+Consequently, combining them would offer limited diversity and might not significantly improve overall performance.
+By contrast, CNN and RNN (correlation: 0.71) exhibit moderate similarity, providing enough redundancy for reliability while maintaining enough difference for enhanced coverage.
+Similarly, GNN and LSTM show the lowest correlation (0.55), suggesting that each model captures distinct aspects of the data and could thus complement one another well.
+Therefore, in our framework, CNN and RNN are aggregated via the DOWA operator, while GNN and LSTM are combined using the IOWA operator.
+This diversity-based selection ensures a balanced trade-off between redundancy (stability) and complementarity (robustness), ultimately improving the ensemble’s ability to detect fraud across a variety of transaction patterns.
+Figure 4: Correlation matrix illustrating the diversity among predictions from RNN, GNN, CNN, and LSTM. Darker shades indicate higher agreement, whereas lighter shades reflect lower correlation.
+II-D2 Attention Layer: DOWA and IOWA Combinations
+In our proposed framework, we first introduce an attention layer that uses the DOWA operator to combine the soft outputs of CNN and RNN.
+DOWA assigns higher weights to classifiers whose predictions lie closer to the overall mean, preventing any single classifier with inconsistent outputs from excessively skewing the final aggregate.
+Since CNN and RNN exhibit a moderate correlation (0.71), this approach balances reliability and diversity by emphasizing their areas of agreement while retaining their distinct predictive signals, ultimately yielding a more robust and stable ensemble.
+Building on this, we extend the attention layer by incorporating the IOWA operator to merge the soft outputs of LSTM and GNN.
+In contrast to DOWA, which highlights predictions near the mean, IOWA weights inputs according to their order and assigned significance, making it especially suitable for combining classifiers with low correlation.
+Given that LSTM and GNN demonstrate the lowest correlation (0.55), this weighted strategy exploits their complementary strengths, LSTM’s ability to capture temporal patterns and GNN’s capacity for graph-structured data representation.
+By carefully assigning weights through IOWA, our framework amplifies reliable signals from each model while minimizing the influence of outlier predictions, further strengthening the ensemble’s overall performance in fraud detection.
+Therefore, the DOWA operator is used to compute weights for the RNN and CNN classifiers, and the IOWA operator does the same for the GNN and LSTM classifiers.
+Once these weights are learned, the soft probabilities from each first-layer classifier are combined accordingly and sorted in descending order.
+For a sample $i\in\{1,2,\dots,N\}$, the DOWA operator produces a two-dimensional output of the form $F_{\mathrm{DOWA}}=\bigl{[}f\bigl{(}p_{\mathrm{RNN}}^{0},p_{\mathrm{CNN}}^{0}\bigr{)},\,f\bigl{(}p_{\mathrm{RNN}}^{1},p_{\mathrm{CNN}}^{1}\bigr{)}\bigr{]}^{T},$ while the IOWA operator yields
+$F_{\mathrm{IOWA}}=\bigl{[}f\bigl{(}p_{\mathrm{GNN}}^{0},p_{\mathrm{LSTM}}^{0}\bigr{)},\,f\bigl{(}p_{\mathrm{GNN}}^{1},p_{\mathrm{LSTM}}^{1}\bigr{)}\bigr{]}^{T}$.
+Here, $p_{\mathrm{RNN}}^{0}$ denotes the probability that sample $i$ belongs to class 0 as predicted by the RNN, and $f\bigl{(}p_{\mathrm{RNN}}^{0},p_{\mathrm{CNN}}^{0}\bigr{)}$ represents the fused value of the respective class-0 probabilities from the RNN and CNN.
+Because each of these operators aggregates the outputs of two classifiers, they effectively map from $\mathbb{R}^{4}$ (the four class probabilities, two from each classifier) to a reduced space $\mathbb{R}^{2}$.
+II-D3 Confidence-Aware Combination (CAC) Layer
+After generating two fused probability vectors, $F_{\mathrm{DOWA}}$ from the CNN–RNN pair and $F_{\mathrm{IOWA}}$ from the GNN–LSTM pair, the framework applies a confidence-based mechanism to decide which set of fused probabilities is passed to the meta-learner.
+Specifically, we quantify uncertainty by comparing the absolute difference between class-0 and class-1 probabilities for each aggregator.
+Formally, for each sample $i$, $\Delta_{\mathrm{DOWA}}=\Bigl{\lvert}f\bigl{(}p_{\mathrm{RNN}}^{0},\,p_{\mathrm{CNN}}^{0}\bigr{)}\;-\;f\bigl{(}p_{\mathrm{RNN}}^{1},\,p_{\mathrm{CNN}}^{1}\bigr{)}\Bigr{\rvert}\quad\text{and}\quad\Delta_{\mathrm{IOWA}}=\Bigl{\lvert}f\bigl{(}p_{\mathrm{GNN}}^{0},\,p_{\mathrm{LSTM}}^{0}\bigr{)}\;-\;f\bigl{(}p_{\mathrm{GNN}}^{1},\,p_{\mathrm{LSTM}}^{1}\bigr{)}\Bigr{\rvert}.$
+We interpret the operator with the smaller absolute difference as having higher uncertainty.
+Consequently, if $\Delta_{\mathrm{IOWA}}\ \leq\Delta_{\mathrm{DOWA}}$, we feed $F_{\mathrm{DOWA}}$ (the DOWA-fused output) into a multi-layer perceptron (MLP) meta-learner; otherwise, we select $F_{\mathrm{IOWA}}$.
+As detailed in Algorithm 1, this decision process is repeated for each sample.
+By always routing the more confident fused probabilities to the meta-learner, the final classification performance is enhanced, particularly under challenging conditions such as imbalanced datasets.
+Algorithm 1 : Confidence-Aware aggregator selection choosing between the DOWA- and IOWA-fused outputs based on uncertainty measures for each sample.
+Input: $f_{i}(p_{RNN}^{0},p_{CNN}^{0})$, $f_{i}(p_{RNN}^{1},p_{CNN}^{1})$, $f_{i}(p_{GNN}^{0},p_{LSTM}^{0})$, $f_{i}(p_{GNN}^{1},p_{LSTM}^{1})$
 Output: One of the aggregate values, $F_{DOWA}$ or $F_{IOWA}$, for each sample
@@
 while $i\leq N$ do
-     if $|f_{i}(p_{BT}^{0},p_{ET}^{0},p_{AB}^{0})-f_{i}(p_{BT}^{1},p_{ET}^{1},p_{AB}^{1})|\geq|f_{i}(p_{SGD}^{0},p_{SVM}^{0},p_{MLP}^{0})-f_{i}(p_{SGD}^{1},p_{SVM}^{1},p_{MLP}^{1})|$ then
+     if $|f_{i}(p_{RNN}^{0},p_{CNN}^{0})-f_{i}(p_{RNN}^{1},p_{CNN}^{1})|\geq|f_{i}(p_{GNN}^{0},p_{LSTM}^{0})-f_{i}(p_{GNN}^{1},p_{LSTM}^{1})|$ then
          return $F_{DOWA}$
@@
 end while
-In fact, for each sample, this layer decides to intelligently select one of the two vectors, $F_{DOWA}$ or $F_{IOWA}$.
-The logic of selection is that whichever of the two operators, DOWA or IOWA, has greater confidence in a sample belonging to one of the two classes, 0 or 1, the aggregated value by that operator is chosen.
-In this way, we use a smart selection of aggregated values by both operators to feed the meta-learner.
-It can be said that the selection layer synergistically combines the results from these two operators.
-III-D4 Meta-learner selection
-The stacking process is significantly influenced by a meta-learner, as it plays a crucial role in determining the classifier’s performance.
-The selection of a classifier for meta learning is crucial as it directly influences the stacked ensemble’s generalization on unseen instances and, subsequently, its overall predictive accuracy.
-In the context of the stacking architecture, the Ridge classifier can be considered a favorable option due to its moderate complexity, which facilitates interpretability.
-Additionally, it incorporates regularization techniques to mitigate overfitting and effectively handles potential multicollinearity among the features used for learning.
-In Table II, the test results are presented, indicating that the Ridge classifier outperforms other base classifiers.
-TABLE I: The degree of accuracy achieved by the first layer classifiers when evaluated on the test data.
-Classifier Accuracy
-Boosted Trees (BT) $93.64\pm 0.22\%$
-Stochastic Gradient Descent (SGD) $92.48\pm 0.13\%$
-Extra Trees (ET) $94.09\pm 0.31\%$
-AdaBoost (AB) $95.33\pm 0.19\%$
-Support vector Machine (SVM) $94.77\pm 0.25\%$
-Multi-Layer Perceptron (MLP) $95.87\pm 0.23\%$
-TABLE II: Accuracy metric for choosing the meta-learner.
-Learner Accuracy
-Decision Tree $95.32\pm 0.16\%$
-Gradient Boosting $96.71\pm 0.13\%$
-K-nearest Neighbor $96.83\pm 0.11\%$
-Logistic Regression $98.44\pm 0.19\%$
-Ridge Classifier $99.95\pm 0.04\%$
-CatBoost $97.63\pm 0.19\%$
-IV Findings and Analysis
-In this segment, we delve deeper into the results and present a range of performance indicators to evaluate the effectiveness of the proposed algorithm.
-These include measures such as accuracy, sensitivity (also known as recall), precision, specificity, Matthew’s Correlation Coefficient (MCC), and the F1-Score.
-IV-A Performance metrics
-IV-A1 Precision
-Precision quantifies the ratio of correct positive predictions to the total number of positive predictions made by a model, highlighting the accuracy of its positive predictions.
-$Precision=\frac{TP}{TP+FP}.$ (19)
-IV-A2 Specificity
-Specificity measures the ratio of accurate negative predictions to the total negative predictions made by a model, underlining the precision of its negative predictions, e.g.,
-$Specificity=\frac{TN}{TN+FP}.$ (20)
-IV-A3 Accuracy
-Accuracy assesses the total proportion of both true positive and true negative predictions made by a model compared to the overall number of predictions, reflecting the general reliability of the model’s predictions, e.g.,
-$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}.$ (21)
-IV-A4 Sensitivity
-Sensitivity quantifies the percentage of true positive instances accurately identified by a model, emphasizing the model’s capacity to detect all positive cases, e.g.,
-$Sensitivity=\frac{TP}{TP+FN}.$ (22)
-IV-A5 MCC
-Unlike other metrics that may only focus on the positive class or are overly optimistic in imbalanced datasets, MCC provides a balanced measure, even if the classes are of very different sizes.
-MCC is defined mathematically as:
-$\small MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}.$ (23)
-IV-A6 F1-Score
-The F1-score is a harmonized measure that considers precision and recall.
-It offers a singular metric that equilibrates the balance between precision and recall, proving valuable in situations involving datasets with imbalances, e.g.,
-$F1-Score=\frac{2\times Precision\times sensitivity}{Precision+sensitivity},$ (24)
-where $TP$ is true positive, $FP$ is false positive, $TN$ is true negative and $FN$ is false negative.
-To prevent data leakage at all stages, including predictions by first layer classifiers, getting weights through a learning process by the IOWA operator, and the final prediction by the meta-learner, the 10-fold cross-validation method was used.
-The exceptional performance of the proposed architecture on the test data aligns with the findings reported in Table III.
+II-D4 Meta-Learner Configuration
+In our final classification stage, we employ a MLP as the meta-learner to process the fused probabilities (i.e., $F_{\mathrm{DOWA}}$ or $F_{\mathrm{IOWA}}$ selected by the confidence-aware combination layer.
+The MLP architecture consists of two hidden layers, each containing $32$ neurons with ReLU activation, to capture non-linear relationships between the fused features.
+A dropout rate of $0.2$ is applied to mitigate overfitting and improve generalization.
+We train the MLP using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of $64$, and early stopping based on a validation loss threshold.
+This configuration ensures both effective learning and a balance between model complexity and computational efficiency.
+Figure 5: Confusion matrix for Dataset 1, using the proposed ensemble model with an 80-20 stratified train-test split.
+Figure 6: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 1.
+Figure 7: Confusion matrix for Dataset 2, using the proposed ensemble model with an 80-20 stratified train-test split.
+Figure 8: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 2.
+Figure 9: Confusion matrix for Dataset 3, where the ensemble model was trained on balanced Dataset 1 and tested on imbalanced Dataset 2.
+Figure 10: Performance comparison of CNN, RNN, LSTM, GNN, and the proposed ensemble model on Dataset 3.
+Figure 11: Impact of incorporating the attention and confidence-aware combination layers in the stacking model on the test data of Dataset 1.
+Figure 12: Impact of incorporating the attention and confidence-aware combination layers in the stacking model on the test data of Dataset 2.
+III Results
+In this section, we delve deeper into the results of our proposed algorithm and present a comprehensive set of performance metrics, including $A_{c}$, recall ($R_{e}$), precision ($P_{r}$), specificity ($S_{p}$), and the F1-Score ($F_{1}$).
+The subsequent figures illustrate how the attention-based ensemble model performs on three different datasets.
+For each dataset, we provide two visualizations: a confusion matrix, which displays classification outcomes for the test set, green areas indicate correct predictions and red areas denote misclassifications, and a performance comparison graph, contrasting the ensemble’s key evaluation metrics against those of the individual deep learning models (CNN, RNN, LSTM, and GNN).
+The fitted probability threshold is shown in each confusion matrix, defining the boundary for classifying instances as either fraudulent (1) or legitimate (0).
+Additionally, a bar chart adjacent to each matrix offers a quick snapshot of how many samples fall into each category.
+This side-by-side comparison underscores the benefits of our attention-based approach in terms of $A_{c}$, $R_{e}$, $P_{r}$, $S_{p}$, and the $F_{1}$.
+Dataset 1 (balanced): In Fig.
+5 and Fig.
+6, the model is trained and tested with an 80–20 stratified split on a balanced dataset (Dataset 1).
+As shown in the confusion matrix (Fig.
+5), the proposed model achieves a high true negative rate, effectively identifying legitimate transactions.
+However, the number of correctly identified fraud cases remains limited, an expected result given the overall lower prevalence of fraud in this dataset.
+The performance comparison (Fig.
+6) demonstrates that the ensemble model consistently surpasses individual deep learning models, with a particularly strong showing in $R_{e}$, a critical metric for fraud detection.
+Dataset 2 (heavily imbalanced): For Dataset 2 (Fig.
+7 and Fig.
+8), the same 80–20 stratified split is applied, but the task becomes more challenging due to a significantly skewed class distribution.
+The confusion matrix (Fig.
+7) reveals an increased number of false negatives, indicating that detecting fraudulent transactions is harder under these conditions.
+This likely stems from distinct fraud patterns and their lower representation in Dataset 2.
+Nonetheless, the ensemble model still outperforms each individual model, as evident in Fig.
+8, maintaining higher $R_{e}$ and $S_{p}$, even under difficult, highly imbalanced circumstances.
+Dataset 3 (cross-dataset evaluation): In Dataset 3 (Fig.
+9 and Fig.
+10), we adopt a different training strategy: Dataset 1 (balanced) is used for training, while Dataset 2 (imbalanced) is reserved for testing.
+The confusion matrix (Fig.
+9) shows that, although the model tends to be conservative, leading to a higher false positive rate, it nonetheless captures a substantial number of fraudulent transactions.
 Fig.
-7 presents the receiver operating characteristic (ROC) curve, which showcases the correlation between the true positive rate (TPR) and the false positive rate (FPR) across different classification thresholds.
-The ROC curves for both classes demonstrate a significant level of overlap, suggesting comparable performance between the two classes.
-The ROC curve’s area under the curve (AUC) functions as a metric to assess classifier performance.
-A higher AUC value, approaching 1, signifies superior classifier performance.
-The suggested model demonstrates the AUC value of 1.
-Figure 7: ROC curve along with corresponding AUC on test data.
-TABLE III: Performance metrics of the proposed method.
-Metric Value
-Precision $100\%$
-Accuracy $99.95\%$
-Sensitivity $99.90\%$
-Specificity $100\%$
-MCC $99.91\%$
-F1-Score $99.94\%$
-V Prospects for Future Studies
+10 confirms that, despite this trade-off, our ensemble model continues to outperform the standalone classifiers, demonstrating a strong balance between $R_{e}$ and $S_{p}$ when transitioning from balanced training data to highly imbalanced real-world scenarios.
+Fig.
+11 and Fig.
+12 evaluate the impact of incorporating attention-based fusion and confidence-aware combination layers into the stacking model by comparing two versions: a baseline model without these layers and an enhanced version that integrates them.
+The results show that adding these extra layers significantly improves performance.
+This improvement indicates that intelligently combining model outputs and reducing uncertainty in final predictions enhances fraud detection effectiveness.
+TABLE II: Performance comparison of various ensemble learning models and the proposed attention-based stacking model for CCF detection across multiple evaluation metrics.
+Datasets LightGBM RF XGBoost DF AdaBoost ET BDT Proposed
+Dataset 1 $P_{r}$ 0.9313 0.8727 0.9280 0.9238 0.9281 0.8853 0.8815 0.9984
+$R_{e}$ 0.8606 0.9172 0.8172 0.8592 0.8595 0.8931 0.9114 1.0000
+$F_{1}$ 0.8945 0.8942 0.8693 0.8901 0.8923 0.8885 0.8959 0.9992
+Dataset 2 $P_{r}$ 0.8847 0.8158 0.8814 0.8799 0.8875 0.7433 0.8375 0.9091
+$R_{e}$ 0.7596 0.8204 0.7400 0.7298 0.7596 0.8827 0.8327 0.8081
+$F_{1}$ 0.8172 0.8181 0.8042 0.7975 0.8162 0.8079 0.8347 0.8556
+Dataset 3 $P_{r}$ 0.0003 0.0012 0.0008 0.0006 0.0018 0.0007 0.0010 0.0095
+$R_{e}$ 0.7155 0.7850 0.7395 0.7104 0.7157 0.7785 0.7874 0.8069
+$F_{1}$ 0.0005 0.0023 0.0015 0.0012 0.0035 0.0014 0.0019 0.0187
+Table II presents a comparative evaluation of various ensemble learning models, including LightGBM, Random Forest (RF), XGBoost, Deep Forest (DF), AdaBoost, Extra Trees (ET), and Boosted Decision Trees (BDT), against the proposed attention-based stacking model for CCF detection.
+The models are assessed across three datasets and multiple evaluation metrics: $P_{r}$, $R_{e}$, and $F_{1}$.
+The proposed model consistently outperforms others, achieving the highest $P_{r}$, $R_{e}$, and $F_{1}$, especially in Dataset 1 ($F_{1}$ = 0.9992) and Dataset 2 ($F_{1}$ = 0.8556).
+Since the features in two datasets were anonymized, we assumed their equivalence when training on Dataset 1 and testing on Dataset 2.
+However, differences in underlying distributions due to potential variations in feature anonymization may have contributed to the lower $F_{1}$ score observed in Dataset 3 ($F_{1}$ = 0.0187).
+This suggests that while the proposed model generalizes well across datasets with similar feature representations, discrepancies introduced by anonymization could affect its ability to transfer knowledge effectively across different datasets.
+Overall, the results across all three datasets confirm the effectiveness and adaptability of the proposed ensemble approach.
+On the balanced dataset (Dataset 1), the model demonstrates high reliability, particularly in correctly identifying legitimate transactions and maintaining a strong $R_{e}$ rate.
+When faced with a heavily imbalanced dataset (Dataset 2), the method preserves its comparative advantage over individual models, underscoring its robustness in challenging real-world conditions.
+Finally, in the cross-dataset evaluation (Dataset 3), the ensemble model again proves its versatility, striking an efficient trade-off between $R_{e}$ and $S_{p}$ when transitioning from a balanced training set to testing on imbalanced data.
+IV Discussion
+Fraud detection, which has gained significant research interest in recent years, has been tackled using a wide range of methodologies and algorithms.
+The following paragraphs highlight notable studies in this area.
+Fanai et al.
+[30] leveraged a deep Autoencoder for representation learning.
+The Autoencoder reduced the dimensionality of the data, extracting compact and informative representations, which were then used to train models, such as DNN, RNN, and a hybrid CNN-RNN.
+The Bayesian optimization algorithm was employed to fine-tune hyperparameters.
+Tian et al.
+[31] employed adaptive sampling and aggregation-based graph neural network (ASA-GNN).
+This model addressed challenges such as noisy data, fraudsters’ camouflage behavior, and the over-smoothing issue in GNNs.
+ASA-GNN employed a neighbor sampling strategy that used cosine similarity to filter irrelevant nodes.
+Zhu et al.
+[32] proposed a noisy-sample-removed under-sampling scheme for imbalanced classification.
+This approach addressed the negative impact of noisy samples on classifier performance.
+The method utilized K-means clustering to remove noisy samples from both majority and minority classes.
+The study [33] proposed an attention-based ensemble model.
+The model improved generalization and reduce variance by introducing a feature-wise attention mechanism and feature diversity regularization.
+The attention mechanism learned the most informative features, while the regularization term optimized the balance between bias and variance.
+The study [34] focus on a federated learning approach which enabled the development of a global model by aggregating locally computed updates from distributed datasets without sharing raw data, ensuring privacy.
+To tackle class imbalance, the study implemented both individual and hybrid resampling techniques, such as synthetic minority over-sampling technique (SMOTE) and adaptive synthetic (ADASYN).
+Zhu et al.
+[35] proposed an adaptive heterogeneous framework integrating deep RL and attention mechanisms.
+It used a reinforcement reward mechanism to select source domain instances and a CNN with an attention module to extract semantic features.
+Tang et al.
+[36] concentrated on a federated graph learning approach, combining federated learning and GNN to model transaction relationships while preserving privacy.
+A graph extension algorithm and adaptive aggregation improved cross-institutional collaboration and addressed data imbalance.
+TABLE III: Comparison summary of the proposed method with existing approaches.
+Work Dataset Objective Deep Learning Model Performance
+2022[6]
+Public: 284 807 samples,
+492 fraudulent cases
+Private: 5 120 000 transactions,
+140 000 fraudulent records
+Binary
+Time-aware
+historical-attention-
+based LSTM
+$P_{r}$ = 0.504
+$R_{e}$ = 0.996
+$F_{1}$ = 0.669
+$P_{r}$ = 0.921
+$R_{e}$ = 0.972
+$F_{1}$ = 0.946
+2023[30]
+Public: 284 807 samples,
+492 fraudulent cases
+Private: 1000 samples,
+30% labled
+as bad credit
+Binary
+Deep neural network
++ RNN
++ Hybric CNN + RNN
+$F_{1}$ = 0.837
+AUC = 0.921
+$F_{1}$ = 0.807
+AUC = 0.908
+2023[31]
+Private: 5 120 000 samples,
+140 000 fraudulent cases
+Public: 160 764 samples,
+(28% fraudulent cases)
+Public: 20 000 samples,
+(balanced dataset)
+Binary
+Aggregation-based
+graph neural
+network
+$R_{e}$ = 0.884
+$F_{1}$ = 0.904
+AUC = 0.922
+$R_{e}$ = 0.737
+$F_{1}$ = 0.774
+AUC = 0.835
+$R_{e}$ = 0.731
+$F_{1}$ = 0.593
+AUC = 0.571
+2023[32]
+Private: 614 samples,
+192 fraudulent cases
+Private: 3 075 samples,
+448 fraudulent cases
+Private: 30 000 samples,
+6 636 fraudulent cases
+Binary
+Decision tree +
+Logistic regression +
+Random forest
+as base classifier
+in the NUS framework
+$F_{1}$ = 0.890
+G-mean = 0.910
+$F_{1}$ = 0.860
+G-mean = 0.870
+$F_{1}$ = 0.830
+G-mean = 0.850
+2023[33]
+Public: 284 807 samples,
+492 fraudulent cases
+Private: 3 500 000 samples,
+fraud rate 1.59%
+Binary
+A feature-wise
+attention-based
+boosting ensemble
+model
+$F_{1}$ = 0.885
+$F_{1}$ = 0.929
+2024[34]
+Public: 284 807 samples,
+492 fraudulent cases
+Binary
+Federated CNN
+models and
+base classifiers
+$A_{c}$ = 0.999
+$P_{r}$ = 0.826
+$R_{e}$ = 0.809
+2024[35]
+Private: 5 125 107 samples,
+147 829 fraudulent cases
+Public: 284 807 samples,
+492 fraudulent cases
+Binary
+A CNN with
+attention
+mechanism
+$P_{r}$ = 0.921
+$F_{1}$ = 0.946
+$P_{r}$ = 0.902
+$F_{1}$ = 0.880
+2024[36]
+Public: 284 807 samples,
+492 fraudulent cases
+590 540 samples
+Binary
+Federated graph
+learning
+$P_{r}$ = 0.925
+$R_{e}$ = 0.910
+$F_{1}$ = 0.918
+$A_{c}$ = 0.935
+$P_{r}$ = 0.710
+$R_{e}$ = 0.615
+$F_{1}$ = 0.640
+$A_{c}$ = 0.800
+2025[37]
+Public: 284 807 samples,
+492 fraudulent cases
+Binary
+Sequential detection:
+LSTM
+Non-Sequential
+detection:
+LightGBM
+$F_{1}$ = 0.850
+AUC = 0.870
+$F_{1}$ = 0.870
+AUC = 0.960
+Our work
+Public: 568 630 samples,
+284 615 fraudulent cases
+Public: 284 807 samples,
+492 fraudulent cases
+Public: 853 437 samples,
+284 807 fraudulent cases
+Binary
+Attention-based
+ensemble of CNNs
+and GNN
+$A_{c}$ = 0.9992
+$F_{1}$ = 0.9992
+Pre = 0.9984
+$R_{e}$ = 1.0
+$A_{c}$= 0.9995
+$F_{1}$ = 0.8556
+Pre = 0.9091
+$R_{e}$ = 0.8081
+$A_{c}$ = 0.830
+$R_{e}$ = 0.8069
+$S_{p}$ = 0.8539
+The study [37] developed integrated resampling methods and data-dependent classifiers.
+It combines one-class SVM with SMOTE and random under-sampling to preserve the distribution of fraud instances.
+The framework outputs are analyzed using LightGBM for non-sequential fraud detection and LSTM for sequential fraud detection.
+Table III compares this study’s approach with other recent fraud detection methods that employ deep neural networks.
+While many existing solutions have achieved respectable performance, our ensemble framework distinguishes itself through its uncertainty-aware design, combining DOWA, IOWA, and a confidence-driven gating mechanism, to effectively handle both balanced and highly imbalanced datasets.
+This architecture not only delivers strong $A_{c}$ and $R_{e}$ but also remains computationally feasible, making it adaptable for deployment in real-time payment gateways.
+Moreover, by integrating SHAP-based feature selection, each chosen attribute can be mapped to concrete transactional insights, thereby enhancing transparency for financial stakeholders.
+Such interpretability is pivotal for fostering trust in AI-driven fraud detection systems, as compliance teams and analysts can readily validate the rationale behind each flagged transaction.
+IV-A Prospects for Future Studies
 In future research, we aim to explore:
-  • In the proposed architecture, we randomly selected 6 classifiers as the first layer classifiers.
+  • In complex tasks, by adding multiple layers of prediction, attention, and selection to the traditional stacking classifier structure, a richer model can identify complex patterns, which could be the focus of future research.
+  • Future studies are recommended to focus on other aggregation methods, such as fuzzy integral operators and approaches with a higher level of complexity than OWA operators.
+  • Credit card transaction patterns can shift rapidly over time due to evolving fraud tactics or changes in consumer behavior.
+Incorporating online or incremental learning approaches could allow the model to adapt continuously and maintain high performance under these dynamic conditions.
+  • Deploying the proposed approach in real-world payment systems requires further examination of computational efficiency.
+Investigations into parallelization, cloud-based infrastructures, or lightweight model variants would help ensure the method can scale for high-throughput, real-time fraud detection scenarios.
+  • We randomly selected four models as the first layer classifiers.
 Future work could optimize the number of first layer classifiers by relying on forward selection and backward elimination approaches.
-  • Future studies are recommended to focus on other aggregation methods, such as fuzzy integral operators and approaches with a higher level of complexity than OWA operators.
-  • There are various criteria for selecting first layer classifiers that future work could employ and compare.
-  • Upcoming studies could implement the suggested architecture for additional datasets that differ in terms of data origin and class balance and evaluate its performance.
-VI Conclusion
-The importance of detecting CCF lies in its ability to protect financial assets, maintain consumer confidence, ensure regulatory compliance, and enhance the overall security and integrity of the financial system.
-In this study, we introduced a novel architecture that is entirely based on integration for the detection of CCF.
-This architecture combines two traditional stacking architectures in parallel through the aid of two key layers including attention and selection.
-The attention layer utilizes two aggregation operators, DOWA and IOWA, to decide how to combine the first layer classifiers’ predictions.
-Each aggregation operator is tasked with combining predictions of three distinct classes.
-The selection layer decides which one of the two operators, DOWA or IOWA, should be considered for feeding the meta-learner for each sample, depending on which operator has more confidence in assigning the sample to one of the two classes, zero or one.
-This architecture, based on combining multiple classifiers, has led to improved overall accuracy, reliability, and decision-making process.
-This technique leverages the strengths of different classifiers and mitigates their individual weaknesses, potentially leading to more robust and accurate predictions, especially in complex or challenging classification tasks.
+V Conclusion
+Given the accelerating shift toward digital transactions and the alarming surge in credit card fraud, effective detection systems have become critically important for financial security.
+This study introduced an attention-based stacking framework for CCF detection, incorporating both DOWA and IOWA aggregation strategies, along with a confidence-aware combination layer, to address the challenges of data imbalance, model uncertainty, and interpretability.
+Empirical evaluation on three datasets, ranging from balanced to heavily imbalanced, highlighted the method’s resilience and superiority over individual classifiers, consistently achieving strong $R_{e}$ and $S_{p}$.
+Additionally, by integrating SHAP to illuminate the most influential features, the model promotes transparency and fosters stakeholder trust.
+Notably, this framework is adaptable to varying computational constraints: in resource-limited settings, users may opt for one of the standalone classifiers to reduce complexity and speed up inference, whereas those with more robust computational resources can deploy the full ensemble approach for enhanced $A_{c}$ and robustness.